OpenStudy (anonymous):

can someone help me? its a long question has 4 parts! will give medal

OpenStudy (anonymous):

Create a rational expression to be your game piece. You may choose from the list of factors below or make your own. There must be a variable term in both the numerator and denominator. (5x) (2x) (x + 4) (x – 5) (2x + 1) (3x + 5)

OpenStudy (owlcoffee):

hmmm So all we hace to do is create a rational expression, huh? okay, I'll make my own just as an example, i'll start with something simple: \[(3x+5)\] I want to express it as something else, so this would be totally "legal" to do: \[(9x+10)-(6x+5)\] This looks very boring and simple, let's add complexity: \[(\sqrt{(9x+10)}+\sqrt{(6x+5)})(\sqrt{(9x+10)}-\sqrt{(6x+5)})\] We know that every number is divided by one so I'll transform that "1" into something else: \[\frac{ (\sqrt{(9x+10)}+\sqrt{(6x+5)})(\sqrt{(9x+10)}-\sqrt{(6x+5)}) }{ \frac{ x ^{2} }{ x ^{2} } }\] Doing the fractionary multiplication: \[\frac{ x ^{2}(\sqrt{(9x+10)}+\sqrt{(6x+5)})(\sqrt{(9x+10)}-\sqrt{(6x+5)}) }{ x ^{2} }\] I'll stop here, it looks perfect! It's just an example, so try creating your own by playing with the variables :)

OpenStudy (anonymous):

could i use that???

OpenStudy (anonymous):

it looks great!

OpenStudy (owlcoffee):

Well, you could try to create your own, but you can sue mine of course. I created it parting from one of the options, try creating yours, I'll help you out.

OpenStudy (anonymous):

theres a second part though Turn one. Flip your coin and perform the appropriate operation. Explain to the game master how to add your rational expression to the one on the correct space. Use complete sentences.

OpenStudy (anonymous):

can you show me using (3x + 5) @Owlcoffee

OpenStudy (owlcoffee):

so they want us to add the rational expression we created to one in the spaces?

OpenStudy (anonymous):

yea i think so

OpenStudy (owlcoffee):

okay: so we have to use common denominator here: our expression added (3x+5). \[\frac{ x ^{2}(\sqrt{(9x+10)}+\sqrt{(6x+5)})(\sqrt{(9x+10)}-\sqrt{(6x+5)}) }{ x ^{2} }+(3x+5)\] in order to add them, we have to remember common denominator wich allows us to sum fractions: \[\frac{ a }{ b }+\frac{ c }{ d }=\frac{ ad+cb }{ bd }\] We'll have to apply that same thing to our problem, where (3x+5) acts like "c" and x^2 as "b", but there's a 1 under (3x+5) that acts like "d" so let's apply it: \[\frac{ x ^{2}(\sqrt{(9x+10)}+\sqrt{(6x+5)})(\sqrt{(9x+10)}-\sqrt{(6x+5)})+(3x+5)x ^{2} }{ x ^{2} }\] and there we go, they are added.

OpenStudy (anonymous):

how would i explain that?

OpenStudy (anonymous):

whats the end result agiain? @Owlcoffee

OpenStudy (anonymous):

and can u delete that last post?

OpenStudy (owlcoffee):

It was that huge rational equality, on the last post.

OpenStudy (anonymous):

this is the other part Turn two. Flip your coin and perform the appropriate operation. Discuss and identify any possible restrictions that exist with (or in) the resulting rational expression.

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

In that part we have to see what value we can't give x. Try looking for any indeterminaton on the rational expression we summed in the last part. (I would show you but I'm in class right now)